10 g of solute A and 20 g of solute B are dissolved in 500 ml water the solution has the same osmotic pressure as 6.67 g of A and 30 g of B are dissolved in the same volume of water at same temperature if the ratio of molar masses of A and B is x/y find x+y

Grade:12

1 Answers

Vikas TU
7 years ago
 
For first soln.,
O.P1 = iCST =>1*((10/x + 20/y)/0.5)*S*T............(1)
For second soln.,
O.P2 = 1*(6.67/x + 30/y)/0.5)*S*T................(2)
Now OP1 = OP2
Therefore,
(10/x + 20/y) = (6.67/x + 30/y)
3.33/x = 10/y
x/y = 3.33/10  = > 333/1000
x+y => 1333